OpenStudy (anonymous):

Solve for x. 5x + 3(x - 2) = -2(x + 1)

OpenStudy (ivancsc1996):

First you apply distributive property on the parenthesis. ¿Can you?

OpenStudy (anonymous):

i know that but, how do you get a number from something that is obviously not balanced

OpenStudy (ivancsc1996):

Yes it is balanced. You just need to do what i told you and solve from there. You get x=4/10

OpenStudy (anonymous):

8x+6=-2x+2, so after that i dont know what to do. i get lost.

OpenStudy (ivancsc1996):

Ok, I see you problem, it is the signs. 6 is negative, since it is 3 times -2, and 2 is negative, since it is -2 times 1. Be careful with signs. It should be: 8x-6=-2x-2. Then you need to put all of the terms with x on one side and the other on the other side. ¿Do you know how to do it?

OpenStudy (anonymous):

i dont know how you would do that though.

OpenStudy (ivancsc1996):

Ok, I'll explain. We need to get all of the x terms on one side and the other on the other side. First we will try to get the "-2x" to the left side of the equation. To do this, we just need to add "-2x" to both sides of the equation. Look what happens:\[8x-6=-2x-2 \rightarrow 8x-6+2x=-2x-2+2x \]So, as you can see, the "-2x" will cancel with the "2x" leaving you with:\[8x-6+2x=-2\]When we added the "2x" on each side, we didn't mess up the equation, since we did it on BOTH sides. Thats the fundamental property of maths. What ever you do on one side, you have to do it on the other side. Now if we add the "8x" and the "2x" we get:\[10x-6=-2\]Now try getting the "-6" term to the right side of the equation, just like I did with the "-2x".